# Observations of microlensed images with dual-field interferometry: on-sky demonstration and prospects

### v1

@@ 3.2 VLTI Data
OGLE-2023-BLG-0061/KMT-2023-BLG-0496 was considered a promising candidate for VLTI observations on 2023 July 20. Although the event was still before the peak, the available data predicted a relatively long timescale ($t_{\rm E}\approx 100$ d). At the same time, the optical brightness indicated that the microlens parallax should be robustly constrained. A nearby ($13.2^{\prime\prime}$) bright ($G=13.7$, $K=9.8$) star 2MASS 17430426-3515195 could serve as a fringe-tracking and adaptive-optics guide star. The first GRAVITY Wide observations were secured with UTs on 2023 July 29 (“epoch 1”). We obtained four sets of medium-resolution observations ($R=\lambda/\Delta\lambda\approx 500$) and one set of low-resolution data ($R\approx 22$), a new mode never tried before with GRAVITY Wide. The first three medium-resolution exposures were of low quality, and the resulting closure phases were very noisy. We, therefore, decided not to use them in the subsequent analysis. The preliminary modeling carried out at that time indicated that the medium- and low-resolution data were inconsistent with each other.
@@
We thus attempted to collect additional VLTI data to study the magnitude of possible systematics affecting the low-resolution data. We triggered on-axis GRAVITY observations with ATs on 2023 August 23, when the event was approaching the peak (at magnification $\sim 90$), but the data could not be collected. The next (successful) attempt to gather additional VLTI UTs observations took place on 2023 September 29 (“epoch 2”), when we secured six low-resolution exposures with GRAVITY Wide. Table 1 presents the log of VLTI observations. Each low-resolution observation consisted of twelve 30 s exposures; the medium-resolution observation consisted of four 100 s exposures. For each night, a bright star pair was observed to center the science fringe with GRAVITY differential delay line before the microlens observation. We did not adjust the science fringe when the telescope was moved to the science target.
@@
The data were reduced with the standard GRAVITY pipeline (version 1.4.2). We first used the Python script run_gravi_reduced.py to reduce the raw data up to applying the pixel-to-visibility matrix (P2VM). The default options were used except that we adopted –gravity_vis.output-phase-sc=SELF_VISPHI to calculate the internal differential phase between each spectral channel and –gravity_vis.opd-pupil-stddev-max-sc=$9$ to ignore the poor pupil measurements in the acquisition camera which do not affect our closure phase measurements. The pipeline performed the bias and sky subtraction, flat fielding, wavelength calibration, and spectral extraction. Application of the P2VM converts the pixel detector counts into complex visibilities, taking into account all instrumental effects, including relative throughput, coherence, phase shift, and cross-talk. The dark, bad pixel, flat field, wavelength calibration, and P2VM matrix data were reduced from the daily calibration data obtained close in time to our observations. We then used run_gravi_trend.py to calibrate the closure phase data. For epoch 1, we used the star pair to center the science fringe and calibrate the medium-resolution data. Unfortunately, we did not have calibrator data observed in the low-resolution data on the same night. For epoch 2, we used a bright star pair observed after the microlens observation for the calibration. The following analyses were based on the closure phase data from the calibrated medium-resolution and uncalibrated low-resolution data from epoch 1, and calibrated low-resolution data from epoch 2. In this way, we used all the data with good quality. Meanwhile, when modeling the low-resolution data from epoch 2, we found that the calibration makes little difference to the closure phase.
4 Light Curve Model
@@
The light curve of the event can be well fitted by a standard point-source point-lens model with the annual parallax effect (which is caused by the orbital motion of the Earth). This model has five free parameters: time of the closest approach between the lens and the source $t_{0}$, their minimum separation (in Einstein radius units) $u_{0}$, the Einstein radius crossing timescale $t_{\rm E}$, and North and East components of the microlensing parallax vector $\boldsymbol{\pi}_{\rm E}=(\pi_{\textrm{E},N},\pi_{\textrm{E},E})$. The latter is a vector quantity whose direction is parallel to the direction of the relative lens–source proper motion $\boldsymbol{\mu}_{\rm rel}$. The magnification is calculated using the formula:
@@
where $\Delta s$ is the projected position of the Sun. Two models with different signs of $u_{0}$ are possible; both models are almost perfectly degenerate, with the $u_{0}<0$ being preferred by only $\Delta\chi^{2}=3.1$. The best-fit model (with $u_{0}>0$) is presented in Figure 2. The best-fit parameters and their uncertainties are reported in Table 2. In Table 2, we also report the best-fit source magnitude $I_{\rm s}$, baseline magnitude $I_{0}$, and dimensionless blending parameter $f_{\rm s}$.
@@
Table 2: Best-fit Parameters of the Light Curve Model
@@
Figure 3 shows the constraints on the microlensing parallax vector derived using data from different observatories. Blue, red, green, and orange contours mark constraints from OGLE, KMTC, KMTA, and KMTS data, respectively, whereas the solid black contours mark the best-fit model to all data. The East component of $\boldsymbol{\pi}_{\rm E}$, which is parallel to the projected acceleration of the Sun, is relatively well measured in all data sets. However, the North component $\pi_{\textrm{E},N}$ (perpendicular to the projected acceleration of the Sun) has a considerably larger uncertainty. The North component is also more susceptible to noise in the data (Gould et al., 1994; Smith et al., 2003; Gould, 2004); hence, slightly different values of $\pi_{\textrm{E},N}$ are determined using different data sets. The best-fit parameters for individual data sets are reported in Tables 3 and 4.
Figure 3: Constraints on the microlensing parallax derived from different photometric data sets (blue – OGLE, red – KMTC, orange – KMTS, green – KMTA) and all light curve data (black contours). The dotted and dashed lines mark constraints from the two epochs of VLTI data, respectively.
Table 3: Best-fit Parameters of the Light Curve Model $(u_{0}>0)$ for Individual Data Sets
@@
Table 4: Best-fit Parameters of the Light Curve Model $(u_{0}<0)$ for Individual Data Sets
@@
The light curve model of a microlensing event makes it possible to predict the brightness ratio $\eta$ between the minor and major image at a given epoch of interferometric observations. Moreover, the detection of the microlensing parallax in the light curve allows us to predict the orientation of microlensed images in the sky. The position angle (north through east) of the microlensing parallax vector can be calculated using formula
@@
The angle $\phi$ between the source–lens relative proper motion and source–lens relative position at a given time can also be calculated directly from the parameters of the light curve model:
@@
Note that in the limit of no parallax, this equation simplifies to $\phi(t)=\arctan(u_{0}t_{\rm E}/(t-t_{0}))$, which is equivalent to Equation (8) derived by Dong et al. (2019). Then, following Dong et al. (2019), the position angle of the minor image relative to the major image (north through east) is simply $\mathrm{PA}=\Phi_{\pi}+\phi$ if $u_{0}>0$ or $\mathrm{PA}=\Phi_{\pi}-\phi$ if $u_{0}<0$. Note that Dong et al. (2019) introduced a different definition of the position angle of the major image relative to the minor image $\psi=\mathrm{PA}+\pi$.
@@
Table 2 presents the predicted lens–source separation $u$ (in Einstein radius units), the flux ratio between the minor and major image $\eta$, and the position angle of the images $\mathrm{PA}$ for both epochs of VLTI observations. Note that the expected lens–source separation and flux ratio of the images are virtually identical for $u_{0}>0$ and $u_{0}<0$ models. However, the expected position angles are different. For the positive $u_{0}$ solution, we can predict $\mathrm{PA}=-164.83\pm 0.95$ deg and $\mathrm{PA}=31.52\pm 0.58$ deg for both VLTI epochs. For the negative $u_{0}$ solution, we predict $\mathrm{PA}=-156.60\pm 0.68$ deg and $\mathrm{PA}=14.29\pm 1.06$ deg, respectively.
@@
We also note that the flux ratio of the images is tightly constrained by the light curve model, with a precision better than 0.3%. Conversely, the uncertainty of the position angle is dominated by the uncertainty in the determination of the angle $\Phi_{\pi}$, which itself is dominated by the uncertainty of $\pi_{\textrm{E},N}$. As we observed small systematic differences between $\pi_{\textrm{E},N}$ determined using different data sets (Figure 3), these differences propagate to systematic variations in $\Phi_{\pi}$ and therefore $\mathrm{PA}$. In particular, the angle $\Phi_{\pi}$ can vary from $18.56$ to $25.90$ deg (Tables 3 and 4).
5 Closure Phase Models
@@
We first separately analyze the VLTI data collected during epochs 1a, 1b, and 2, because they were taken at different times and using different instrument configurations. That will allow us to study the consistency between the model parameters derived using different data sets and the consistency with the light curve model (Sections 5.1, 5.2, and 5.3). In Section 5.4, we combine all VLTI data sets to derive the final parameters of the system. The results presented in this section were independently checked using the PMOIRED software [^5] https://github.com/amerand/PMOIRED (Mérand, 2022) and we found virtually identical results.
5.1 Binary-star Model
@@
We start by fitting a simple binary-star model to the closure phase data. We assume that both images of the source star can be considered as pointlike. We place the major image in the origin of the coordinate system, whereas the position of the minor image is parameterized by a vector $(\Delta\alpha,\Delta\delta)$. If we denote the flux ratio between the minor and major image as $\eta$, then the complex visibility is
$\mathrm{VIS}=\frac{1+\eta\exp\left(-\frac{2\pi i}{\lambda}\left(U\Delta\alpha+V\Delta\delta\right)\right)}{1+\eta},$ (6)
where $\lambda$ is the wavelength of observations, and $(U,V)$ is the separation of the telescopes in the UV plane. The visibility is calculated for each pair of telescopes, and then the closure phase T3 is calculated for each triangle. Four possible closure phases can be calculated for VLTI; in theory, one of them is not independent. However, since the measured closure phases may be affected by noise in the data, we employ all four closure phase sets in the fits.
We find the best-fit parameters by maximizing the log-likelihood function defined as:
@@
where $n_{\rm exp}$ is the number of exposures in the epoch and $\Lambda$ is the number of spectral channels. Since the error bars calculated from the pipeline may be underestimated, we add a constant error term $\sigma_{0}$ in quadrature. The best-fit parameters and their uncertainties are calculated using the Markov chain Monte Carlo (MCMC) algorithm coded by Foreman-Mackey et al. (2013). We assume flat (non-informative) priors on all parameters of the model.
Table 5: Best-fit Parameters of the Binary-star Closure Phase Model
@@
$\chi^{2}$/d.o.f. $824.5/924$ $49.6/48$ $282.6/308$
Figure 4: Closure phase data for the VLTI epoch 1. Color solid lines mark the best-fit binary-star model.
Figure 5: Closure phase data for the VLTI epoch 2. Color solid lines mark the best-fit binary-star model.
Results of the fits, separately for VLTI epochs 1a, 1b, and 2, are reported in Table 5. The best-fitting closure phase models are shown in Figures 4 and 5. In addition to four model parameters, we report four derived quantities: separation between the images $s=\sqrt{\Delta\alpha^{2}+\Delta\delta^{2}}$, their position angle (north through east) $\mathrm{PA}=\arctan\Delta\alpha/\Delta\delta$, source–lens separation $u$ (in Einstein radius units), and the angular Einstein radius $\theta_{\rm E}$. The latter quantity is calculated using the formula
@@
Results of the fits to individual epoch 2 exposures are presented in Table 6.
@@
Further checks are possible because the flux ratio $\eta$ and position angle $\mathrm{PA}$ of the images can be independently measured using the light curve data (Section 4). In particular, there are two possible light curve models differing by the sign of $u_{0}$, which predict the position angles of the microlensed images of $(-164.8^{\circ},31.5^{\circ})$ (positive $u_{0}$) or $(-156.6^{\circ},14.3^{\circ})$ (negative $u_{0}$), during the VLTI epochs 1 and 2, respectively. The measured angles ($-162.9^{\circ},26.7^{\circ}$) (Table 5) seem to favor the $u_{0}>0$ model.
However, while the light curve and closure phase models agree well for epoch 1, the expected position angles differ by almost $4.8^{\circ}$ (that is, $7.5\sigma$) during epoch 2. That is illustrated in Figure 6, which presents $1,2,3\sigma$ confidence ellipses in the $(\mathrm{PA},\eta)$ plane. Gray contours mark the constraints from three different VLTI data sets, whereas black contours were calculated based on the light curve model ($u_{0}>0$). Blue, red, green, and orange contours mark constraints on $(\mathrm{PA},\eta)$ from OGLE, KMTC, KMTA, and KMTS data, respectively.
The position angle of the images measured from interferometric observations can be projected onto the $(\pi_{\textrm{E},N},\pi_{\textrm{E},E})$ plane (Figure 3). The dotted and dashed lines in this figure correspond to the best-fitting values of PA during epochs 1a and 2, respectively. The dashed line does not intersect color contours from the light curve model, which further exemplifies the tension discussed above.
[…]
Figure 6: Constraints on the flux ratio and position angle of the microlensed images measured from the VLTI data (gray contours). The color contours mark the constrains from the individual photometric data sets (blue – OGLE, red – KMTC, orange – KMTS, green – KMTA), the black contours are measured using all photometric data combined.
Table 6: Best-fit Parameters of the Binary-star Closure Phase Model for Individual Epoch 2 Data
@@
$\chi^{2}$/d.o.f. $51.6/48$ $50.6/48$ $47.5/48$ $40.2/48$ $28.6/48$ $46.5/48$
5.2 Luminous-lens Model
Despite the slight tensions discussed above, the binary-star model overall describes the closure phase data well (Figures 4 and 5). We now modify the model to include a possibility of blended light coming from the lens itself or a luminous companion (either to the lens or source). We changed the primary parameters of the model to the lens–source separation $u$, the position angle of the lens relative to the source $\mathrm{PA}$ (which is equal to the position angle of the minor image relative to the major image), and the angular Einstein radius $\theta_{\rm E}$, which provide a more natural description of a microlensing event than the binary-star model parameters. We keep $\sigma_{0}$ fixed at 6.990, 2.593, and 2.509 deg for epoch 1a, 1b, and 2, respectively.

### v2

@@ 3.2 VLTI Data
OGLE-2023-BLG-0061/KMT-2023-BLG-0496 was considered a promising candidate for VLTI observations on 2023 July 20. Although the event was still before the peak, the available data predicted a relatively long timescale ($t_{\rm E}\approx 100$ days). At the same time, the optical brightness indicated that the microlens parallax should be robustly constrained. A nearby ($13.2^{\prime\prime}$) bright ($G=13.7$, $K=9.8$) star, 2MASS 17430426-3515195, could serve as a fringe-tracking and adaptive optics guide star. The first GRAVITY Wide observations were secured with UTs on 2023 July 29 (“epoch 1”). We obtained four sets of medium-resolution observations ($R=\lambda/\Delta\lambda\approx 500$) and one set of low-resolution data ($R\approx 22$), a new mode never tried before with GRAVITY Wide. The first three medium-resolution exposures were of low quality, and the resulting closure phases were very noisy. We therefore decided not to use them in the subsequent analysis. The preliminary modeling carried out at that time indicated that the medium- and low-resolution data were inconsistent with each other.
@@
We thus attempted to collect additional VLTI data to study the magnitude of the possible systematics affecting the low-resolution data. We triggered on-axis GRAVITY observations with ATs on 2023 August 23, when the event was approaching the peak (at magnification $\sim 90$), but the data could not be collected. The next (successful) attempt to gather additional VLTI UT observations took place on 2023 September 29 (“epoch 2”), when we secured six low-resolution exposures with GRAVITY Wide. Table 1 presents the log of VLTI observations and Figure 2 shows the coverage of the UV plane during both VLTI epochs. Each low-resolution observation consisted of twelve 30 s exposures; the medium-resolution observation consisted of four 100 s exposures. For each night, a bright-star pair was observed to center the science fringe with the GRAVITY differential delay line before the microlens observation. We did not adjust the science fringe when the telescope was moved to the science target.
@@
The data were reduced with the standard GRAVITY pipeline (version 1.4.2). We first used the Python script run_gravi_reduced.py to reduce the raw data and apply the pixel-to-visibility matrix (P2VM). The default options were used, except that we adopted –gravity_vis.output-phase-sc=SELF_VISPHI to calculate the internal differential phase between each spectral channel and –gravity_vis.opd-pupil-stddev-max-sc=$9$ to ignore the poor pupil measurements in the acquisition camera, which do not affect our closure-phase measurements. The pipeline performed the bias and sky subtraction, flat-fielding, wavelength calibration, and spectral extraction. The application of the P2VM converts the pixel detector counts into complex visibilities, taking into account all instrumental effects, including relative throughput, coherence, phase shift, and crosstalk. The dark, bad-pixel, flat-field, wavelength calibration, and P2VM matrix data were reduced from the daily calibration data obtained close in time to our observations. We then used run_gravi_trend.py to calibrate the closure-phase data. For epoch 1, we used the star pair to center the science fringe and calibrate the medium-resolution data. Unfortunately, we did not have calibrator data observed in the low-resolution data on the same night. For epoch 2, we used a bright-star pair observed after the microlens observation for the calibration. The following analyses are based on the closure-phase data from the calibrated medium-resolution and uncalibrated low-resolution data from epoch 1 and the calibrated low-resolution data from epoch 2. In this way, we use all the data with good quality. Meanwhile, when modeling the low-resolution data from epoch 2, we found that the calibration makes little difference to the closure phase.
4 Light-curve Model
@@
The light curve of the event can be well fitted by a standard point-source point-lens model with the annual parallax effect (which is caused by the orbital motion of the Earth). This model has five free parameters: the time of closest approach between the lens and the source $t_{0}$, their minimum separation (in Einstein radius units) $u_{0}$, the Einstein radius crossing timescale $t_{\rm E}$, and the northern and eastern components of the microlensing parallax vector $\boldsymbol{\pi}_{\rm E}=(\pi_{\textrm{E},N},\pi_{\textrm{E},E})$. The latter is a vector quantity whose direction is parallel to the direction of the relative lens–source proper motion $\boldsymbol{\mu}_{\rm rel}$. The magnification is calculated using the formula
@@
where $\Delta s$ is the projected position of the Sun. Two models with different signs of $u_{0}$ are possible; both models are almost perfectly degenerate, with the $u_{0}<0$ model being preferred by only $\Delta\chi^{2}=3.1$. The best-fit model (with $u_{0}>0$) is presented in Figure 3. The best-fit parameters and their uncertainties are reported in Table 2. In Table 2, we also report the best-fit source magnitude $I_{\rm s}$, baseline magnitude $I_{0}$, and dimensionless blending parameter $f_{\rm s}$.
@@
Table 2: Best-fit Parameters of the Light-curve Model
@@
Note. — Parameters denoted by (1) and (2) were calculated for the VLTI epochs 1 and 2, respectively.
Figure 4 shows the constraints on the microlensing parallax vector derived using data from different observatories. The blue, red, green, and orange contours mark the constraints from OGLE, KMTC, KMTA, and KMTS data, respectively, whereas the solid black contours mark the best-fit model to all data. The eastern component of $\boldsymbol{\pi}_{\rm E}$, which is parallel to the projected acceleration of the Sun, is relatively well measured in all data sets. However, the northern component $\pi_{\textrm{E},N}$ (perpendicular to the projected acceleration of the Sun) has a considerably larger uncertainty. The northern component is also more susceptible to noise in the data (Gould et al., 1994; Smith et al., 2003; Gould, 2004); hence, slightly different values of $\pi_{\textrm{E},N}$ are determined using different data sets. The best-fit parameters for individual data sets are reported in Tables 3 and 4.
Figure 4: Constraints on the microlensing parallax derived from different photometric data sets (blue—OGLE; red—KMTC; orange—KMTS; and green—KMTA) and all light-curve data (black contours). The dotted and dashed lines mark the constraints from the two epochs of VLTI data, respectively.
Table 3: Best-fit Parameters of the Light-curve Model $(u_{0}>0)$ for Individual Data Sets
@@
Table 4: Best-fit Parameters of the Light-curve Model $(u_{0}<0)$ for Individual Data Sets
@@
The light-curve model of a microlensing event makes it possible to predict the brightness ratio $\eta$ between the minor and major image at a given epoch of interferometric observations. Moreover, the detection of the microlensing parallax in the light curve allows us to predict the orientation of microlensed images in the sky. The position angle (PA; north through east) of the microlensing parallax vector can be calculated using the formula
@@
The angle $\phi$ between the source–lens relative proper motion and source–lens relative position at a given time can also be calculated directly from the parameters of the light-curve model:
@@
Note that in the limit of no parallax, this equation simplifies to $\phi(t)=\arctan(u_{0}t_{\rm E}/(t-t_{0}))$, which is equivalent to Equation (8) derived by Dong et al. (2019). Then, following Dong et al. (2019), the position angle of the minor image relative to the major image (north through east) is simply $\mathrm{PA}=\Phi_{\pi}+\phi$ if $u_{0}>0$ or $\mathrm{PA}=\Phi_{\pi}-\phi$ if $u_{0}<0$. These two cases depend on the trajectory of the lens relative to the source: $u_{0}>0$ if the lens passes the source on its right and $u_{0}<0$ on its left (Skowron et al., 2011). Note that Dong et al. (2019) introduced a different definition of the position angle of the major image relative to the minor image: $\psi=\mathrm{PA}+\pi$.
@@
Table 2 presents the predicted lens–source separation $u$ (in Einstein radius units), the flux ratio between the minor and major image $\eta$, and the position angle of the images $\mathrm{PA}$ for both epochs of VLTI observations. Note that the expected lens–source separation and flux ratio of the images are virtually identical for the $u_{0}>0$ and $u_{0}<0$ models. However, the expected position angles are different. For the positive $u_{0}$ solution, we can predict $\mathrm{PA}=-164.83\pm 0.95$ deg and $\mathrm{PA}=31.52\pm 0.58$ deg for both VLTI epochs. For the negative $u_{0}$ solution, we predict $\mathrm{PA}=-156.60\pm 0.68$ deg and $\mathrm{PA}=14.29\pm 1.06$ deg, respectively.
@@
We also note that the flux ratio of the images is tightly constrained by the light-curve model, with a precision better than 0.3%. Conversely, the uncertainty of the position angle is dominated by the uncertainty in the determination of the angle $\Phi_{\pi}$, which itself is dominated by the uncertainty of $\pi_{\textrm{E},N}$. As we observed small systematic differences between $\pi_{\textrm{E},N}$ determined using different data sets (Figure 4), these differences propagate to systematic variations in $\Phi_{\pi}$ and therefore $\mathrm{PA}$. In particular, the angle $\Phi_{\pi}$ can vary from $18.56$ to $25.90$ deg (Tables 3 and 4).
5 Closure-phase Models
@@
The VLTI/GRAVITY observations of OGLE-2023-BLG-0061/KMT-2023-BLG-0496 provide a complementary view of the microlensing event. In particular, the spatial resolution of the interferometer allows us to resolve the microlensed images, providing a precise measurement of the angular Einstein radius. This information is not included in the light curve of the event. We first separately analyze the VLTI data collected during epochs 1a, 1b, and 2, because they were taken at different times and using different instrument configurations. That will allow us to study the consistency between the model parameters derived using different data sets and the consistency with the light-curve model (Sections 5.1, 5.2, and 5.3). In Section 5.4, we combine all VLTI data sets to derive the final parameters of the system. The results presented in this section were independently checked using the PMOIRED software [^6] https://github.com/amerand/PMOIRED (Mérand, 2022) and we found virtually identical results.
5.1 No-lens-light Model
@@
In the simplest case, if the lens is dark and there are no other blended stars in the GRAVITY field of view, the two images of the source created by microlensing look like a mundane “binary star.” We start by fitting such a simple binary star model to the closure-phase data. We assume that both images of the source star can be considered as point-like. We place the major image in the origin of the coordinate system, whereas the position of the minor image is parameterized by a vector $(\Delta\alpha,\Delta\delta)$. If we denote the flux ratio between the minor and major image as $\eta$, then the complex visibility is
$\hat{V}=\frac{1+\eta\exp\left(-\frac{2\pi i}{\lambda}\left(U\Delta\alpha+V\Delta\delta\right)\right)}{1+\eta},$ (6)
[…]
We first conduct the grid search to find the initial parameters of the model. We keep the flux ratio fixed to $\eta=0.591$ (epoch 1) and $\eta=0.426$ (epoch 2)—that is, the best-fit values from the light-curve model (Table 2). We then evaluate the goodness-of-the-fit statistic $\chi^{2}$ on a grid of $201\times 201$ points uniformly distributed over the range $-10\,\textrm{mas}\leq(\Delta\alpha,\Delta\delta)\leq 10$ mas. For each case, we find there is only one significant minimum of the $\chi^{2}$ surface.
We then refine the results of the grid search by allowing all parameters to vary. This is achieved by maximizing the log-likelihood function defined as
@@
where $n_{\rm exp}$ is the number of exposures in the epoch and $\Lambda$ is the number of spectral channels. Because the error bars calculated from the pipeline may be underestimated, we add a constant error term $\sigma_{0}$ in quadrature. The best-fit parameters and their uncertainties are calculated using the Markov Chain Monte Carlo (MCMC) algorithm coded by Foreman-Mackey et al. (2013). We assume flat (noninformative) priors on all parameters of the model.
Table 5: Best-fit Parameters of the Closure-phase Model Without the Lens Light
@@
$\chi^{2}$/dof $824.5/924$ $49.6/48$ $282.6/308$
Figure 5: Closure-phase data for the VLTI epoch 1. The colored solid lines mark the best-fit microlensing model without the lens light.
Figure 6: Closure-phase data for the VLTI epoch 2. The colored solid lines mark the best-fit microlensing model without the lens light.
The results of the fits—separately for VLTI epochs 1a, 1b, and 2—are reported in Table 5. The best-fitting closure-phase models are shown in Figures 5 and 6. In addition to four model parameters, we report four derived quantities: the separation between the images $s=\sqrt{\Delta\alpha^{2}+\Delta\delta^{2}}$, their position angle (north through east) $\mathrm{PA}=\arctan\Delta\alpha/\Delta\delta$, the source–lens separation $u$ (in Einstein radius units), and the angular Einstein radius $\theta_{\rm E}$. The latter quantity is calculated using the formula
@@
The results of the fits to the individual epoch 2 exposures are presented in Table 6.
@@
Further checks are possible because the flux ratio $\eta$ and position angle $\mathrm{PA}$ of the images can be independently measured using the light-curve data (Section 4). In particular, there are two possible light-curve models differing by the sign of $u_{0}$, which predict the position angles of the microlensed images as $(-164.8^{\circ},31.5^{\circ})$ (positive $u_{0}$) or $(-156.6^{\circ},14.3^{\circ})$ (negative $u_{0}$), during the VLTI epochs 1 and 2, respectively. The measured angles ($-162.9^{\circ},26.7^{\circ}$) (Table 5) seem to favor the $u_{0}>0$ model.
However, while the light-curve and closure-phase models agree well for epoch 1, the expected position angles differ by almost $4.8^{\circ}$ (that is, $7.5\sigma$) during epoch 2. That is illustrated in Figure 7, which presents $1\sigma$, $2\sigma$, and $3\sigma$ confidence ellipses in the $(\mathrm{PA},\eta)$ plane. The gray contours mark the constraints from three different VLTI data sets, whereas the black contours are calculated based on the light-curve model ($u_{0}>0$). The blue, red, green, and orange contours mark the constraints on $(\mathrm{PA},\eta)$ from OGLE, KMTC, KMTA, and KMTS data, respectively.
The position angle of the images measured from interferometric observations can be projected onto the $(\pi_{\textrm{E},N},\pi_{\textrm{E},E})$ plane (Figure 4). The dotted and dashed lines in this figure correspond to the best-fitting values of the PAs during epochs 1a and 2, respectively. The dashed line does not intersect color contours from the light-curve model, which further exemplifies the tension discussed above.
The light-curve model also predicts the flux ratios between the microlensed images: $\eta=0.5907\pm 0009$ for epoch 1 and $\eta=0.4264\pm 0.0013$ for epoch 2. The former value is in $2.7\sigma\mbox{--}3.5\sigma$ tension with the closure-phase fits to epoch 1a and 1b data, respectively.
Figure 7: Constraints on the flux ratio and position angle of the microlensed images measured from the VLTI data (gray contours). The colored contours mark the constraints from the individual photometric data sets (blue—OGLE; red—KMTC; orange—KMTS; and green—KMTA), while the black contours are measured using all photometric data combined.
Table 6: Best-fit Parameters of the Closure-phase Model for Individual Epoch 2 Data Without the Lens Light
@@
$\chi^{2}$/dof $51.6/48$ $50.6/48$ $47.5/48$ $40.2/48$ $28.6/48$ $46.5/48$
5.2 Luminous Lens Model
Despite the slight tensions discussed above, the simple binary star model overall describes the closure-phase data well (Figures 5 and 6). We now modify the model to include the possibility of blended light coming from the lens itself or a luminous companion (either to the lens or source). We change the primary parameters of the model to the lens–source separation $u$, the position angle of the lens relative to the source $\mathrm{PA}$ (which is equal to the position angle of the minor image relative to the major image), and the angular Einstein radius $\theta_{\rm E}$, which provide a more natural description of a microlensing event than the binary star model parameters. We keep $\sigma_{0}$ fixed at 6.990, 2.593, and 2.509 deg for epochs 1a, 1b, and 2, respectively.